OpenStudy (supernova_sonntag):

At what points, if any, does the circle (x-3)^2+(y-3)^2=25 intersect the line y=2x-1? Use exact answers, do not round. Help??

OpenStudy (anonymous):

\[\left( x-3 \right)^2+\left( 2x-1-3 \right)^2=25\] \[\left( x-3 \right)^2+\left( 2x-4 \right)^2=25\] simplify and find x